### 23 Mar 2026 v2

# On single-frequency asymptotics for

Abstract
We consider damped driven Maxwell–Bloch equations which are finite-dimensional approximation of the damped driven Maxwell–Schrödinger equations. The equations describe a single-mode Maxwell field coupled to a two-level molecule. Our main result is the construction of solutions with single-frequency asymptotics of the Maxwell field in the case of quasiperiodic pumping. The asymptotics hold for solutions with harmonic initial values which are stationary states of averaged equations in the interaction picture.
We calculate all harmonic states and analyse their stability. The calculations rely on the Bloch–Feynman gyroscopic representation of von Neumann equation for the density matrix. The asymptotics follow by application of the averaging theory of the Bogolyubov type. The key role in the application of the averaging theory is played by a special a priori estimate.
MSC classification: 37J40, 58D19, 37J06, 70H33, 34C25, 34C29, 78A40, 78A60.
Keywords: Maxwell–Bloch equations; Bloch–Feynman vector equation; Hamiltonian structure; density matrix; von Neumann equation; pumping; averaging theory; single-frequency asymptotics; quantum optics; laser.
1 Introduction
The Maxwell–Bloch equations (MBE) were introduced by Lamb [25] for the semiclassical description of the laser action [1, 6, 17, 31, 32, 33, 34]. The equations are the Galerkin approximation of the Maxwell–Schrödinger system [5, 16, 19, 21, 22, 27, 29] Our main goal is construction of solutions with single-frequency asymptotics. The asymptotics seem to correspond to the laser coherent radiation, which remains a key mystery of laser action since its discovery around 1960. The damped driven MBE for mixed states read
$\left\{\begin{array}[]{rcl}\dot{A}(t)\!\!&\!\!=\!\!&\!\!B(t),\quad\dot{B}(t)=-\Omega^{2}A(t)\!-\!\gamma B(t)+cj(t)\\ \\ i\hbar\dot{\rho}(t)\!\!&\!\!=\!\!&\!\![H(t),\rho(t)]\end{array}\right|,\qquad j(t)={2}\varkappa\,{\rm Im{\hskip 1.42262pt}}\rho_{21}(t).$ (1.1)
[…]
$H(t)=\left(\!\!\!\begin{array}[]{cc}\hbar\omega_{1}&ia(t)\\ \\ -ia(t)&\hbar\omega_{2}\end{array}\!\!\!\right);\qquad\rho=\left(\!\!\!\begin{array}[]{cc}\rho_{11}&\overline{\rho}_{21}\\ \\ \rho_{21}&\rho_{22}\end{array}\!\!\!\right),\qquad\mathop{\rm tr\,}\nolimits\rho=1,\quad\rho\geq 0,$ (1.2)
where $\hbar\omega_{2}>\hbar\omega_{1}$ are the energy levels of active molecules, and the function $a(t)$ is given by
$a(t):=\frac{\varkappa}{c}[A(t)+A^{e}(t)],\qquad\varkappa=p\,{{\omega}},\quad{\omega}=\omega_{2}-\omega_{1}>0,$ (1.3)
[…]
For solutions to the MBE, the conservation ${\rm tr{\hskip 1.42262pt}}\rho(t)=\mathop{\rm const}\nolimits$ holds since the trace of commutators vanish. We will consider solutions with $\mathop{\rm const}\nolimits=1$; see (1.2). We use the Heaviside–Lorentz units and recall the introduction of the equations in Appendix B.
In Appendix A, we prove the a priori bound for solutions $X(t)=(A(t),B(t),\rho(t))$ to (1.1) with $|p|/\gamma=r$ satisfy the following a priori bounds:
$|X(t)|\leq D_{r}(|X(0)|),\qquad t\geq 0,$ (1.5)
which is proved in Appendix A. The bound implies the well-posedness of the MBE.
Our main goal is asymptotics of the Maxwell amplitudes $A(t)$ and $B(t)$ for solutions $X(t)$ to the MBE as $|p|,\gamma\to 0$. Define complex Maxwell amplitudes by $M(t)=A(t)+iB(t)/\Omega$. Then the first two equations of the MBE are equivalent to
$\dot{M}(t)=-i\Omega M(t)-i\gamma M_{2}(t)+2icp\omega\,{\rm Im{\hskip 1.42262pt}}\rho_{21}/\Omega,\quad{\rm where}\quad M_{2}(t)={\rm Im{\hskip 1.42262pt}}M(t).$ (1.6)
Note that the parameters $|p|,\gamma$ are very small for many types of lasers, see Appendix C. For $p=\gamma=0$, all solutions to (1.6) are single-frequency: $M(t)=e^{-i\Omega t}M(0)$. For small $p$ and $\gamma$, equation (1.6) implies that $M(t)=e^{-i\Omega t}M(0)+\displaystyle\int_{0}^{t}R(s)ds,$ where $\sup\limits_{s\geq 0}|R(s)|={\cal O}(|p|+\gamma)$ by (1.5). Hence, $M(t)=e^{-i\Omega t}M(0)+{\cal O}(|p|t)$ for $p/\gamma=r$ with arbitrary $r\neq 0$. In particular, for solutions with any fixed initial states $M(0)$,
$\max_{t\in[0,|p|^{-1/2}]}|M(t)-e^{-i\Omega t}M(0)|={\cal O}({|p|^{1/2}}),\qquad{p\to 0},\quad{p}/\gamma=r.\quad\qquad\qquad\qquad\qquad$ (1.7)
Our main results show that in the resonance case, when $\Omega=\omega$, the time scale $p^{-1/2}$ in this asymptotics can be extended to $|p|^{-1}$ and even more in the case of special harmonic states of the MBE.
We will represent the density matrix via the Pauli matrices as
[…]
$\dot{\bf X}(t)=p{\bf F}_{r}({\bf X}(t)),\qquad t\geq 0;\qquad r=p/\gamma.$ (1.9)
We define the harmonic states ${\bf X}\in{\mathbb X}$ of the MBE as stationary solutions of (1.9).
We calculate all harmonic states ${\bf X}^{r}=({\bf M}^{r},{\bf S}^{r})\in{\mathbb X}$ of the MBE with $p/\gamma=r$ and show that the set is a union $Z^{r}=Z^{r}_{1}\cup Z^{r}_{2}$ of two smooth 1D manifolds. We linearise the dynamics (1.9) at the harmonic states and calculate the spectra of the linearisations. The calculations show that for $\Omega=\omega$ and $c|r|>|{\bf A}^{e}|$, there exist a nonempty submanifold $Z^{r}_{+}\subset Z^{r}$ which is attractive under the averaged dynamics (1.9). Denote the matrix
$\qquad\qquad\qquad\qquad V_{\Omega}=\left(\begin{array}[]{ccc}0&\Omega&0\\ -\Omega&0&0\\ 0&0&0\end{array}\right).$ (1.10)
Our main results are the following asymptotics for solutions $X(t)=(M(t),S(t))$ to equations (1.1) in the representation (1.8), with a fixed quotient $p/\gamma=r$. Everywhere below we consider the case $p>0$ only since the extension to $p<0$ is obvious.
Theorem $1.$
Let $\Omega=\omega$ and the pumping be quasiperiodic. Then for any $r>0$, the following asymptotics hold.
i) Let the initial state $X(0)=({\bf M},{\bf S})\in Z^{r}$. Then the corresponding solutions admit the following adiabatic asymptotics:
$\max_{t\in[0,p^{-1}]}\Big[|M(t)-e^{-i\Omega t}{\bf M}+|S(t)-e^{V_{\Omega}t}{\bf S}|\Big]={\cal O}(p^{1/2}),\qquad p\to 0.\qquad\qquad\qquad\quad$ (1.11)
ii) Let $cr>|{\bf A}^{e}|$, and $D^{r}_{d}$ denote a suitable subset of the tubular $d$-neighborhood of the stable submanifold $Z^{r}_{+}$ with sufficiently small $d>0$. Then for initial states $X(0)\in D^{r}_{d}$, the corresponding solutions admit the following asymptotics uniformly in initial values $X(0)\in D^{r}_{d}$:
$\max_{t\in[0,p^{-1}]}\Big[|M(t)-e^{-i\Omega t}{\bf M}_{*}|+|S(t)-e^{V_{\Omega}t}{\bf S}_{*}(t)|\Big]={\cal O}[{p^{1/2}}+d],\qquad p\to 0,\qquad$ (1.12)
where $({\bf M}_{*},{\bf S}_{*}(t))\in Z^{r}_{+}$ and ${\bf M}_{*}=-{\bf A}^{e}$ does not depend on $r$ and on $X(0)\in D^{r}_{d}$. The Lebesgue measure
$|D^{r}_{d}|\sim d^{4},\qquad d\to 0.$ (1.13)
Remark 1.2.
Without the constraint $p/\gamma=r$, the a priori bounds (1.5) and the asymptotics (1.11)–(1.12) do not hold. For example, the bounds must be different for small and big $|p|$ for a given dissipation $\gamma>0$. The asymptotics do not hold without this restriction since the limiting amplitudes ${\bf M}^{r}$ depend on $r$.
The asymptotics (1.11) for $S(t)$ reads
$S(t)=e^{V_{\Omega}t}{\bf S}+{\cal O}(p^{1/2}),\qquad t\in[0,p^{-1}]$ (1.14)
which is approximately the precession of $S(t)$ about axis $S_{3}$.
[…]
The calculation of harmonic states and of spectra of the linearisations rely on the Bloch–Feynman representation [7, 14] of the von Neumann equation via the vector $S(t)$ from (1.8). We prove that the harmonic states $({\bf M},{\bf S})$ with ${\bf M}\neq 0$ exist only in the resonance case $\Omega=\omega$, and only if ${\bf A}^{e}\neq 0$. In particular, for the single-frequency pumping $A^{e}(t)={\rm Re{\hskip 1.42262pt}}[{\bf A}_{p}e^{-i\omega_{p}t}]$, the harmonic states with ${\bf M}\neq 0$ exist only in the case ${\bf A}_{p}\neq 0$ and triple resonance
$\Omega=\omega=\omega_{p}.$ (1.15)
The asymptotics (1.11) follows for solutions with every initial state $(M(0),S(0))\in Z^{r}$ by the averaging theory of Bogolyubov type [30, Theorem 4.3.6]. To prove asymptotics (1.12), we combine the theorem with the stability of the branch $Z^{r}_{+}$. Our calculations show that for all harmonic states $({\bf M},{\bf S})\in Z^{r}_{+}$, the component ${\bf M}=-{\bf A}^{e}$. This is why the limiting amplitude $\bf M$$_{*}$ in (1.12) does not depend on $r>0$ and on $X(0)\in D^{r}_{d}$. The key role in the application of the averaging theory is played by the special a priori estimate (1.5).
Remark 1.3.
Our approach relies on the averaging theory which neglects oscillating terms. So, it gives a justification of the “rotating wave approximation”, which is widely used in Quantum Optics [1, 17, 31, 32, 33, 34]. The asymptotics (1.11), (1.12) specify the time scale and the error of such approximations.
Let us comment on related results. The problem of existence of time-periodic solutions to the MBE has been discussed since 1960s. The first results in this direction were obtained recently in [11] and [36] for various versions of the MBE. In [36], the N-th order time-periodic solutions were constructed by perturbation techniques. For the phenomenological model [2, 3], time-periodic solutions were constructed in [11] in the absence of time-periodic pumping for small interaction constants. The solutions are obtained as the result of a bifurcation relying on homotopy invariance of the degree [9] and developing the averaging arguments [10]. The period is determined by bifurcation.
In [20], we have established the existence of solutions with $T$-periodic Maxwell amplitude for any $T$-periodic pumping without smallness conditions.
In [24], we have constructed solutions with asymptotics of type (1.11), (1.12) for the MBE with pure states, using the reduction by the symmetry gauge group. In this case, for any $r>0$, the set of stationary states is discrete and consists of one or two points. The construction of the asymptotics for system (1.1) in the present paper, required completely new technique based on the Bloch–Feynman representation.
Up to our knowledge, the single-frequency asymptotics for the MBE with mixed states were not constructed till now.
Let us comment on our exposition. In Section 2, we construct the representation of the Bloch–Feynman type for the von Neumann equation from MBE. The dynamics on the interaction picture is calculated in Section 3, its averaging in Section 4, and all stationary states of the averaged equations are calculated in Section 5. In Section 6 we analyse the stability of the stationary states, and in Section 8 we prove the single-frequency asymptotics (1.11)–(1.12). In Appendix A we establish the bound (1.5). In Appendix B we comment on the introduction of the MBE with mixed states, and in Appendix C we discuss possible treatment of the laser threshold and laser amplification relying on our results.
Acknowledgements. The authors thank S. Kuksin, M.I. Petelin, A. Shnirelman and H. Spohn for longterm fruitful discussions, and the Institute of Mathematics of BOKU University for the support and hospitality. The research is supported by Austrian Science Fund (FWF) PAT 3476224.
2 The Bloch-Feynman “gyroscopic” representation
Here we represent von Neumann equation from MBE as a “gyroscopic equation” (2.5) applying the Bloch, Feynman & al approach [6, 7, 14, 15]. Hermitian density matrices $\rho$ with $\mathop{\rm tr\,}\nolimits\rho=1$ admit the representation
[…]
Hence, using the Pauli matrices and the nonnegativity $\det\rho=1-S_{1}^{2}-S_{2}^{2}-S_{3}^{2}\geq 0$, we obtain (1.8):
$\rho=\frac{1}{2}\left(\!\!\begin{array}[]{cc}1&0\\ 0&1\end{array}\!\!\right)+\frac{S_{1}}{2}\left(\!\!\begin{array}[]{cc}0&1\\ 1&0\end{array}\!\!\right)+\frac{S_{2}}{2}\left(\!\!\begin{array}[]{cc}0&-i\\ i&0\end{array}\!\!\right)+\frac{S_{3}}{2}\left(\!\!\begin{array}[]{cc}1&0\\ 0&-1\end{array}\!\!\right)=\frac{1}{2}[E+S\cdot\sigma].$ (2.2)
In particular, we can expand the “Hamiltonian operator” (1.2) as
$H(t)=\frac{\hbar}{2}(\omega_{1}+\omega_{2})E+\frac{\hbar}{2}(\omega_{1}-\omega_{2})\sigma_{3}-a(t)\sigma_{2}=\frac{\hbar}{2}(\omega_{1}+\omega_{2})E+\frac{\hbar}{2}(\omega_{1}-\omega_{2})e_{3}\cdot\sigma-a(t)e_{2}\cdot\sigma,$ (2.3)
where $e_{3}=(0,0,1)$ and $e_{2}=(0,1,0)$. Now the von Neumann equation in MBE for $\rho(t)=\frac{1}{2}[E+S(t)\cdot\sigma]$ reads as
$i\hbar\dot{S}(t)\cdot\sigma=2i\frac{\hbar}{2}(\omega_{1}-\omega_{2})[e_{3}\wedge S(t)]\cdot\sigma-a(t)[e_{2}\wedge S(t)]\cdot\sigma=2i\Big[\big(\frac{\hbar}{2}(\omega_{1}-\omega_{2})e_{3}-a(t)e_{2}\big)\wedge S(t)\Big]\cdot\sigma,$ (2.4)
[…]
$\dot{S}(t)=\theta(t)\wedge S(t),\qquad\theta(t)=-\omega\,e_{3}-2\frac{a(t)}{\hbar}e_{2},\quad\omega=\omega_{2}-\omega_{1}.$ (2.5)
The equation can be also written as
$\dot{S}(t)=\Theta(t)S(t),\qquad\Theta(t)=\begin{pmatrix}0&-\theta_{3}(t)&\theta_{2}(t)\\ \theta_{3}(t)&0&-\theta_{1}(t)\\ -\theta_{2}(t)&\theta_{1}(t)&0\end{pmatrix}=\begin{pmatrix}0&\omega&-2a(t)/\hbar\\ -\omega&0&0\\ 2a(t)/\hbar&0&0\end{pmatrix}\in so(3),$ (2.6)
which corresponds to the cross-product representation of the Lie algebra $so(3)$. In particular, the following conservation law holds,
$|S(t)|=\mathop{\rm const}\nolimits,\qquad t\geq 0.$ (2.7)
The current $j(t)=\varkappa S_{2}(t)$, so the MBE (1.1) reduces to (cf. (1.6))
$\left\{\begin{array}[]{rcl}\dot{M}(t)\!\!&\!\!=\!\!&\!\!-i\Omega M(t)-i\gamma M_{2}(t)+ic\varkappa S_{2}(t)/\Omega\\ \\ \dot{S}(t)\!\!&\!\!=\!\!&\!\!{\Theta(t)}S(t)\end{array}\right|,\qquad M(t)=A(t)+iB(t)/\Omega.$ (2.8)
3 Dynamics in the interaction picture
By (2.6), for small $p,\gamma>0$, the system (2.8) is a small perturbation of the unperturbed one,
$\dot{M}(t)=-i\Omega M(t),\qquad\dot{S}(t)=V_{\omega}S(t),$ (3.1)
where $V_{\omega}$ is the matrix (1.10) with $\Omega=\omega$ which coincides with (2.6) in the case $a(t)=0$. Solutions to this system are given by
[…]
Our goal is construction of similar solutions to the perturbed system (2.8),
$M(t)=e^{-i\Omega t}{\mathbb M}(t),\qquad S(t)=e^{V_{\omega}t}{\mathbb S}(t)$ (3.3)
with slowly varying enveloping amplitudes: for a wide interval of time $[0,T(p)]$
$\sup_{t\in[0,T(p)]}\big[|{\mathbb M}(t)-{\mathbb M}(0)|+|{\mathbb S}(t)-{\mathbb S}(0)|\big]\to 0,\qquad p\to 0,\quad p/\gamma=r\neq 0.$ (3.4)
Substituting (3.3) into (2.8), we obtain the dynamical equations for the enveloping amplitudes (the “interaction picture”):
$\left\{\begin{array}[]{rl}\dot{\mathbb M}(t)&=-ie^{i\Omega t}\big[\gamma{\rm Im{\hskip 1.42262pt}}(e^{-i\Omega t}{\mathbb M}(t))-\tilde{\varkappa}(e^{V_{\omega}t}{\mathbb S}(t))_{2}\big]\\ \\ \dot{\mathbb S}(t)&=e^{-V_{\omega}t}({\Theta(t)}-V_{\omega})e^{V_{\omega}t}{\mathbb S}(t)\end{array}\right|,\qquad\tilde{\varkappa}=\frac{c\varkappa}{\Omega}.$ (3.5)
The equations are called as the interaction picture of (2.8). By (2.5), (2.6), and (1.3),
$\Theta(t)-V_{\omega}=\frac{2a(t)}{\hbar}\left(\begin{array}[]{ccc}0&0&-1\\ 0&0&0\\ 1&0&0\end{array}\right),\qquad a(t):=\frac{\varkappa}{c}[A(t)+A^{e}(t)],\qquad\varkappa=p\,{{\omega}}.$ (3.6)
Note that $e^{V_{\omega}t}$ is he dynamical group of the gyroscopic equation (2.5) with the angular velocity $\theta(t)=-\omega e_{3}$. Hence, $e^{V_{\omega}t}$ is the rotation about $e_{3}$ with the angular velocity $\omega$:
$e^{V_{\omega}t}=\left(\begin{array}[]{ccc}\,\,\,\,\,\,\,\,\cos\omega t&\sin\omega t&0\\ -\sin\omega t&\cos\omega t&0\\ 0&0&1\end{array}\right)$ (3.7)
Substituting (3.6) and (3.7) into (3.5), we obtain
$\left\{\begin{array}[]{rl}\dot{\mathbb M}(t)=\!\!&\!\!-pie^{i\Omega t}\!\big[\gamma_{1}({\mathbb M}_{2}(t)\cos\Omega t-{\mathbb M}_{1}(t)\sin\Omega t)-\varkappa_{1}(-{\mathbb S}_{1}(t)\sin\omega t+{\mathbb S}_{2}(t)\cos\omega t)\big],\\ \\ \dot{\mathbb S}(t)=\!\!&\!\!-pb[A(t)\!+\!A^{e}(t)]\big[{\mathbb S}_{3}(t)(e_{1}\cos\omega t+e_{2}\sin\omega t)-e_{3}({\mathbb S}_{1}(t)\cos\omega t+{\mathbb S}_{2}(t)\sin\omega t)\big],\end{array}\right$.
where
$\gamma_{1}=\frac{\gamma}{p}=\frac{1}{r},\qquad\varkappa_{1}=\frac{\tilde{\varkappa}}{p}=\frac{c\,{\omega}}{\Omega},\qquad b=\displaystyle\frac{2\omega}{c\hbar}.$ (3.8)
The equations can be written as
$\dot{\mathbb M}(t)=p\,f_{r}({\mathbb M}(t),{\mathbb S}(t),t),\qquad\dot{\mathbb S}(t)=p\,g_{r}({\mathbb M}(t),{\mathbb S}(t),t),$ (3.9)
where the functions $f$ and $g$ are given by
$\left\{\!\!\!\!\!\begin{array}[]{rl}f_{r}({\mathbb M},{\mathbb S},t)\!\!=\!\!\!&\!\!\!\!\!\big[\!-\!i\cos\Omega t\!+\!\sin\Omega t\big]\big[\gamma_{1}({\mathbb M}_{2}\cos\Omega t\!-\!{\mathbb M}_{1}\sin\Omega t)\!-\!\varkappa_{1}(\!-{\mathbb S}_{1}\sin\omega t\!+\!{\mathbb S}_{2}\cos\omega t)\big]\\ \\ g_{r}({\mathbb M},{\mathbb S},t)\!\!=\!\!\!&\!\!\!\!\!-b\big[{\mathbb M}_{1}\cos\Omega t\!+\!{\mathbb M}_{2}\sin\Omega t\!+\!A^{e}(t)\big]\big[{\mathbb S}_{3}(e_{1}\cos\omega t\!+\!e_{2}\sin\omega t)\!-\!e_{3}({\mathbb S}_{1}\cos\omega t\!+\!{\mathbb S}_{2}\sin\omega t)\!\big]\end{array}\!\!\!\!\right|\!.\!\!$ (3.10)
Remark 3.1.
It is important that the coefficients $\gamma_{1},\varkappa_{1},b$ depend only on $\Omega,\omega,r$. Hence, for any fixed $\Omega,\omega>0$ and $r\neq 0$, the asymptotics of solutions to systems(3.9) as $p\to 0$ and $p/\gamma=r$ can be calculated by methods of the averaging theory [8, 30].
4 The averaging
The averaged equations (3.9) read
$\dot{\bf M}(t)=p\,\overline{f}_{r}({\bf M}(t),{\bf S}(t)),\qquad\dot{\bf S}(t)=p\,\overline{g}_{r}({\bf M}(t),{\bf S}(t)),$ (4.1)
where
$\overline{f}_{r}({\bf M},{\bf S})=\langle f_{r}({\bf M},{\bf S},\cdot)\rangle=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}f_{r}({\bf M},{\bf S},t)dt,\qquad\overline{g}_{r}({\bf M},{\bf S})=\langle g_{r}({\bf M},{\bf S},\cdot)\rangle.$ (4.2)
Let us calculate the averages (4.2). The results differ drastically for the resonance case $\Omega=\omega$ and non-resonance $\Omega\neq\omega$. In notation (1.4), ${\bf A}^{e}={\bf A}^{e}_{1}+i{\bf A}^{e}_{2}$, where
${\bf A}^{e}=2\langle A^{e}(t)e^{i\Omega t}\rangle,\qquad{\bf A}^{e}_{1}=2\langle A^{e}(t)\cos\Omega t\rangle,\qquad{\bf A}^{e}_{2}=2\langle A^{e}(t)\sin\Omega t\rangle.$ (4.3)
Resonance case $\Omega=\omega$. Using the expressions (3.10), we obtain the averaged vector field
$\overline{f}_{r}({\bf M},{\bf S})=-\frac{i}{2}\![\gamma_{1}{\bf M}_{2}-\varkappa_{1}{\bf S}_{2}\big]-\frac{1}{2}\big[\gamma_{1}{\bf M}_{1}-\varkappa_{1}{\bf S}_{1}]=-\frac{\gamma_{1}}{2}[{\bf M}_{1}+i{\bf M}_{2}]+\frac{\varkappa_{1}}{2}[{\bf S}_{1}+i{\bf S}_{2}].$ (4.4)
$\overline{g}_{r}({\bf M},{\bf S})$ $=$ $-\frac{b}{2}\big[{\bf M}_{1}[-e_{3}{\bf S}_{1}+{\bf S}_{3}e_{1}]+{\bf M}_{2}\![-e_{3}{\bf S}_{2}+{\bf S}_{3}e_{2}]-\!e_{3}({\bf S}_{1}{\bf A}^{e}_{1}+{\bf S}_{2}{\bf A}^{e}_{2})+\!{\bf S}_{3}(e_{1}{\bf A}^{e}_{1}\!+\!e_{2}{\bf A}^{e}_{2})\big]$ (4.5)
$=$ $-\frac{b}{2}\big[{\bf S}_{3}[e_{1}({\bf M}_{1}+{\bf A}^{e}_{1})+e_{2}({\bf M}_{2}+{\bf A}^{e}_{2})]-e_{3}[{\bf S}_{1}({\bf M}_{1}+{\bf A}^{e}_{1})+{\bf S}_{2}({\bf M}_{2}+{\bf A}^{e}_{2})]\big].$
Non-resonance case $\Omega\neq\omega$. In this case, the calculations (4.4) simplify to
$\overline{f}_{r}({\bf M},{\bf S})=\displaystyle-\frac{\gamma_{1}}{2}{\bf M},$ (4.6)
which implies nonexistence of stationary states with nonzero Maxwell field for the system (4.1), and also the decay ${\bf M}(t)=\displaystyle{\bf M}(0)e^{-\frac{\gamma}{2}t}$. In contrast, the resonance equation (4.4) includes the interaction term which can prevent the decay of the Maxwell amplitude, that is expected physically.
The averaged system. By (4.4) and (4.5), the averaged equations (4.1) (or (1.9)) in the resonance case read as
$\left\{\begin{array}[]{rcl}\dot{\bf M}(t)&=&-\frac{p}{2}\big[\gamma_{1}[{\bf M}_{1}+i{\bf M}_{2}]-\varkappa_{1}[{\bf S}_{1}+i{\bf S}_{2}]\big]\\ \\ \dot{\bf S}(t)&=&-\frac{pb}{2}\big[{\bf S}_{3}[e_{1}({\bf M}_{1}+{\bf A}^{e}_{1})+e_{2}({\bf M}_{2}+{\bf A}^{e}_{2})]-e_{3}[{\bf S}_{1}({\bf M}_{1}+{\bf A}^{e}_{1})+{\bf S}_{2}({\bf M}_{2}+{\bf A}^{e}_{2})]\big]\end{array}\right|.$ (4.7)
5 Stationary states for the averaged dynamics (harmonic states)
In this section, we calculate all harmonic states $({\bf M},{\bf S})$, i.e., stationary states for the averaged equations (4.7), in the resonance case $\Omega=\omega$. The states satisfy
$\left\{\begin{array}[]{rl}0=&\displaystyle\gamma_{1}[{\bf M}_{1}+i{\bf M}_{2}]-\varkappa_{1}[{\bf S}_{1}+i{\bf S}_{2}]\\ \\ 0=&{\bf S}_{3}[e_{1}({\bf M}_{1}+{\bf A}^{e}_{1})+e_{2}({\bf M}_{2}+{\bf A}^{e}_{2})]-e_{3}[{\bf S}_{1}({\bf M}_{1}+{\bf A}^{e}_{1})+{\bf S}_{2}({\bf M}_{2}+{\bf A}^{e}_{2})]\end{array}\right|.$ (5.1)
It is important that the stationary equations depend on $r$ but do not depend on pp. The first equation of (5.1) gives
${\bf S}_{1}=\alpha_{r}{\bf M}_{1},\quad{\bf S}_{2}=\alpha_{r}{\bf M}_{2},\quad{\rm where}\quad\alpha_{r}=\frac{\gamma_{1}}{\varkappa_{1}}=\frac{1}{cr}>0.$ (5.2)
The second equation of (5.1) is equivalent to the system
$\left\{\begin{array}[]{rcl}0&=&{\bf S}_{3}({\bf M}_{1}+{\bf A}^{e}_{1})\\ 0&=&{\bf S}_{3}({\bf M}_{2}+{\bf A}^{e}_{2})\\ 0&=&{\bf S}_{1}({\bf M}_{1}+{\bf A}^{e}_{1})+{\bf S}_{2}({\bf M}_{2}+{\bf A}^{e}_{2})\end{array}\right|.$ (5.3)
The last equation together with (5.2) give ${\bf M}_{1}({\bf M}_{1}+{\bf A}^{e}_{1})+{\bf M}_{2}({\bf M}_{2}+{\bf A}^{e}_{2})=0$ which defines the circle $S:|{\bf M}|^{2}+{\bf M}\cdot{\bf A}^{e}=0$, or equivalently, $|{\bf M}+{\bf A}^{e}/2|^{2}=|{\bf A}^{e}|^{2}/4$, so
${\bf M}={\bf M}(\theta)=-\frac{{\bf A}^{e}}{2}+\frac{|{\bf A}^{e}|}{2}e^{i\theta},\qquad\theta\in[0,2\pi].$ (5.4)
Recall that $|{\bf S}|\leq 1$ by (1.8). Hence, the set of all stationary states with a fixed $r=p/\gamma$ is the union of two real 1D manifolds $Z^{r}=Z^{r}_{1}\cup Z^{r}_{2}$, where
[…]
The second summand $Z^{r}_{2}$ is
[…]
Remark 5.1.
i) $Z^{r}_{1}\neq\emptyset$ for all $r\neq 0$ since ${\bf M}(0)=0$.
ii) $Z^{r}_{2}\neq\emptyset$ iff $|\alpha_{r}{\bf A}^{e}|\leq 1$, i.e. $cr\geq|{\bf A}^{e}|$.
iii) The intersection of the manifolds is at most one point:
$Z^{r}_{1}\cap Z^{r}_{2}=\left\{\begin{array}[]{ll}(-{\bf A}^{e},-\alpha_{r}{\bf A}^{e},0),&\alpha_{r}|{\bf A}^{e}|\leq 1\\ \emptyset,&\alpha_{r}|{\bf A}^{e}|>1\end{array}\right|.$ (5.7)
Remark 5.2.
[…]
6 Spectra of linearised equations at the harmonic states
Here, we calculate i) linearisations of the averaged system (4.7) at the harmonic states and ii) their spectra. By (4.4) and (4.5), we obtain, omitting the index $r$:
$\left\{\begin{array}[]{rl}\overline{f}_{i}({\bf M},{\bf S})=&-\frac{1}{2}\big[\gamma_{1}{\bf M}_{i}-\varkappa_{1}{\bf S}_{i}\big],\quad i=1,2,\\ \\ \overline{g}_{k}({\bf M},{\bf S})=&-\frac{b}{2}\big[{\bf S}_{3}({\bf M}_{k}+{\bf A}^{e}_{k})\big],\quad k=1,2,\\ \\ \overline{g}_{3}({\bf M},{\bf S})=&\frac{b}{2}\big[{\bf S}_{1}({\bf M}_{1}+{\bf A}^{e}_{1})+{\bf S}_{2}({\bf M}_{2}+{\bf A}^{e}_{2})]\end{array}\right$.
Differentiating, we get for $i,j,k,l=1,2$
$\left\{\!\!\!\!\begin{array}[]{rclrclrcl}\displaystyle\frac{\partial\overline{f}_{i}}{\partial{\bf M}_{j}}&=&-\frac{\gamma_{1}}{2}\delta_{ij},&\displaystyle\frac{\partial\overline{f}_{i}}{\partial{\bf S}_{l}}&=&\frac{\varkappa_{1}}{2}\delta_{il},&\displaystyle\frac{\partial\overline{f}_{i}}{\partial{\bf S}_{3}}&=&0\\ \\ \displaystyle\frac{\partial\overline{g}_{k}}{\partial{\bf M}_{j}}&=&-\frac{b}{2}{\bf S}_{3}\delta_{kj},&\displaystyle\frac{\partial\overline{g}_{k}}{\partial{\bf S}_{l}}&=&0,&\displaystyle\frac{\partial\overline{g}_{k}}{\partial{\bf S}_{3}}&=&-\!\frac{b}{2}({\bf M}_{k}+{\bf A}^{e}_{k})\\ \\ \displaystyle\frac{\partial\overline{g}_{3}}{\partial{\bf M}_{j}}&=&\frac{b}{2}{\bf S}_{j},&\displaystyle\frac{\partial\overline{g}_{3}}{\partial{\bf S}_{l}}&=&\frac{b}{2}({\bf M}_{l}+{\bf A}^{e}_{l}),&\displaystyle\frac{\partial\overline{g}_{3}}{\partial{\bf S}_{3}}&=&0\end{array}\right$.
Thus, we obtain the Jacobian
$J=J({\bf M},{\bf S})=\frac{1}{2}\begin{pmatrix}-\gamma_{1}&0&\varkappa_{1}&0&0\\ 0&-\gamma_{1}&0&\varkappa_{1}&0\\ -b{\bf S}_{3}&0&0&0&-b({\bf M}_{1}+{\bf A}^{e}_{1})\\ 0&-b{\bf S}_{3}&0&0&-b({\bf M}_{2}+{\bf A}^{e}_{2})\\ b{\bf S}_{1}&b{\bf S}_{2}&b({\bf M}_{1}+{\bf A}^{e}_{1})&b({\bf M}_{2}+{\bf A}^{e}_{2})&0\end{pmatrix}$.
Let us consider the cases $({\bf M},{\bf S})\in Z^{r}_{1}$ and $({\bf M},{\bf S})\in Z^{r}_{2}$ separately.
I. For $({\bf M},{\bf S})\in Z^{r}_{1}$, the Jacobian simplifies to
$J=\frac{1}{2}\begin{pmatrix}-{\gamma_{1}}&0&{\varkappa_{1}}&0&0\\ 0&-{\gamma_{1}}&0&{\varkappa_{1}}&0\\ 0&0&0&0&-b{\bf B}_{1}\\ 0&0&0&0&-b{\bf B}_{2}\\ b{\bf S}_{1}&b{\bf S}_{2}&b{\bf B}_{1}&b{\bf B}_{2}&0\end{pmatrix}$,
where we denote ${\bf B}_{k}={\bf M}_{k}+{\bf A}^{e}_{k}$. Hence,
$\det(2J-\lambda)=\begin{bmatrix}-\gamma_{1}-\lambda&0&\varkappa_{1}&0&0\\ 0&-\gamma_{1}-\lambda&0&\varkappa_{1}&0\\ 0&0&-\lambda&0&-b{\bf B}_{1}\\ 0&0&0&-\lambda&-b{\bf B}_{2}\\ b{\bf S}_{1}&b{\bf S}_{2}&b{\bf B}_{1}&b{\bf B}_{2}&-\lambda\end{bmatrix}$
$(\gamma_{1}+\lambda)\begin{bmatrix}-\gamma_{1}-\lambda&0&\varkappa_{1}&0\\ 0&-\lambda&0&-b{\bf B}_{1}\\ 0&0&-\lambda&-b{\bf B}_{2}\\ b{\bf S}_{2}&b{\bf B}_{1}&b{\bf B}_{2}&-\lambda\end{bmatrix}+\varkappa_{1}\begin{bmatrix}0&-\gamma_{1}-\lambda&\varkappa_{1}&0\\ 0&0&0&-b{\bf B}_{1}\\ 0&0&-\lambda&-b{\bf B}_{2}\\ b{\bf S}_{1}&b{\bf S}_{2}&b{\bf B}_{2}&-\lambda\end{bmatrix}$
$(\gamma_{1}+\lambda)^{2}\begin{bmatrix}-\lambda&0&-b{\bf B}_{1}\\ 0&-\lambda&-b{\bf B}_{2}\\ b{\bf B}_{1}&b{\bf B}_{2}&-\lambda\end{bmatrix}-\varkappa_{1}(\gamma_{1}+\lambda)\!\begin{bmatrix}0&-\lambda&-b{\bf B}_{1}\\ 0&0&-b{\bf B}_{2}\\ b{\bf S}_{2}&b{\bf B}_{1}&-\lambda\end{bmatrix}\!+\!\varkappa_{1}b{\bf S}_{1}\!\begin{bmatrix}\gamma_{1}+\lambda&\varkappa_{1}&0\\ 0&0&-b{\bf B}_{1}\\ 0&-\lambda&-b{\bf B}_{2}\end{bmatrix}$
$(\gamma_{1}+\lambda)^{2}\big(-\lambda^{3}-\lambda b^{2}{\bf B}_{1}^{2}-\lambda b^{2}{\bf B}_{2}^{2}\big)-\varkappa_{1}(\gamma_{1}+\lambda)\lambda b^{2}{\bf S}_{2}{\bf B}_{2}-\varkappa_{1}(\gamma_{1}+\lambda)\lambda b^{2}{\bf S}_{1}{\bf B}_{1}$
$(\gamma_{1}+\lambda)\lambda\big[(\gamma_{1}+\lambda)\big(\lambda^{2}+b^{2}({\bf B}_{1}^{2}+{\bf B}_{2}^{2})\big)+\varkappa_{1}b^{2}({\bf S}_{1}{\bf B}_{1}+{\bf S}_{2}{\bf B}_{2})\big]=-(\gamma_{1}+\lambda)^{2}\lambda\big(\lambda^{2}+b^{2}({\bf B}_{1}^{2}+{\bf B}_{2}^{2})\big)$
since ${\bf S}_{1}{\bf B}_{1}+{\bf S}_{2}{\bf B}_{2}=0$ by (5.3). Hence,
$\lambda_{1}=\lambda_{2}=-\gamma_{1},\quad\lambda_{3}=0,\quad\lambda_{4,5}=\pm ib\sqrt{{\bf B}_{1}^{2}+{\bf B}_{2}^{2}}.$
II. For $({\bf M},{\bf S})\in Z^{r}_{2}$, the Jacobian reads
$J=\frac{1}{2}\begin{pmatrix}-\gamma_{1}&0&\varkappa_{1}&0&0\\ 0&-\gamma_{1}&0&\varkappa_{1}&0\\ -b{\bf S}_{3}&0&0&0&0\\ 0&-b{\bf S}_{3}&0&0&0\\ b{\bf S}_{1}&b{\bf S}_{2}&0&0&0\end{pmatrix}$. (6.1)
Hence,
$\det(2J-\lambda)=\begin{bmatrix}-\gamma_{1}-\lambda&0&\varkappa_{1}&0&0\\ 0&-\gamma_{1}-\lambda&0&\varkappa_{1}&0\\ -b{\bf S}_{3}&0&-\lambda&0&0\\ 0&-b{\bf S}_{3}&0&-\lambda&0\\ b{\bf S}_{1}&b{\bf S}_{2}&0&0&-\lambda\end{bmatrix}=-\lambda\begin{bmatrix}-\gamma_{1}-\lambda&0&\varkappa_{1}&0\\ 0&-\gamma_{1}-\lambda&0&\varkappa_{1}\\ -b{\bf S}_{3}&0&-\lambda&0\\ 0&-b{\bf S}_{3}&0&-\lambda\end{bmatrix}$
$\lambda\big[(\gamma_{1}+\lambda)\begin{bmatrix}\gamma_{1}+\lambda&0&\varkappa_{1}\\ 0&-\lambda&0\\ b{\bf S}_{3}&0&-\lambda\end{bmatrix}+b{\bf S}_{3}\begin{bmatrix}0&\varkappa_{1}&0\\ \gamma_{1}+\lambda&0&\varkappa_{1}\\ b{\bf S}_{3}&0&-\lambda\end{bmatrix}\big]$
$\lambda\big[(\gamma_{1}+\lambda)(\lambda^{2}(\gamma_{1}+\lambda)+\lambda\varkappa_{1}b{\bf S}_{3})+b{\bf S}_{3}(\varkappa_{1}^{2}b{\bf S}_{3}+\lambda\varkappa_{1}(\gamma_{1}+\lambda)\big]$
$\lambda\big[\lambda^{2}(\gamma_{1}+\lambda)^{2}+2\lambda(\gamma_{1}+\lambda)b\varkappa_{1}{\bf S}_{3}+b^{2}\varkappa_{1}^{2}{\bf S}_{3}^{2}\big]=-\lambda\big[\lambda^{2}+\lambda\gamma_{1}+b\varkappa_{1}{\bf S}_{3}\big]^{2}.$
Hence, we have the following roots:
$\lambda_{1,2}({\bf S}_{3})=-\frac{\gamma_{1}}{2}+\frac{\sqrt{\gamma_{1}^{2}-4b\varkappa_{1}{\bf S}_{3}}}{2},\qquad\lambda_{3,4}({\bf S}_{3})=-\frac{\gamma_{1}}{2}-\frac{\sqrt{\gamma_{1}^{2}-4b\varkappa_{1}{\bf S}_{3}}}{2},\qquad\lambda_{5}=0.$ (6.2)
By (3.8), the spectrum of the linearised system (4.1) at the stationary states $({\bf M},{\bf S})\in Z^{r}_{2}$ consists of
$p\lambda_{1,2}({\bf S}_{3})=-\frac{\gamma}{2}+\frac{\sqrt{\gamma^{2}-4p^{2}b\varkappa_{1}{\bf S}_{3}}}{2},\qquad p\lambda_{3,4}({\bf S}_{3})=-\frac{\gamma}{2}-\frac{\sqrt{\gamma^{2}-4p^{2}b\varkappa_{1}{\bf S}_{3}}}{2},\qquad\lambda_{5}=0.$
By (3.8), we have $b\varkappa_{1}=\frac{2\omega}{\hbar}>0$. Hence, for $p\neq 0$,
${\rm Re{\hskip 1.42262pt}}\lambda_{k}({\bf S}_{3})<0\qquad{\rm for}\quad k=1,\dots,4\quad{\rm iff}\quad{\bf S}_{3}>0.$ (6.3)
7 Attraction to stable submanifold
Let us assume that $r$ is sufficiently large, so that
$cr>|{\bf A}^{e}|.$ (7.1)
Then $Z_{2}^{r}\neq\emptyset$ and $|\alpha_{r}{\bf A}^{e}|=|{\bf A}^{e}|/(cr)<1$, and we can rewrite (5.6) as
$Z^{r}_{2}=\{{\bf Z}^{r}_{2}({\bf S}_{3})=(-{\bf A}^{e},-\alpha_{r}{\bf A}^{e}_{1},-\alpha_{r}{\bf A}^{e}_{2},{\bf S}_{3}):{\bf S}_{3}\in[-\beta_{r},\beta_{r}]\},\qquad\beta_{r}:=\sqrt{1-\alpha_{r}^{2}|{\bf A}^{e}|^{2}}>0.$ (7.2)
For $0\leq a\leq b\leq 1$ and $d>0$ denote
$Z^{r}_{+}(a,b)=\{{\bf Z}^{r}_{2}({\bf S}_{3}):{\bf S}_{3}\in[a,b]\},\qquad{\cal U}^{r}_{d}(a,b)=\{{\bf Z}^{r}_{2}({\bf S}_{3})+{\bf N}:{\bf N}\bot TZ^{r}_{2}({\bf S}_{3}),\,\,\,{\bf S}_{3}\in[a,b],\,\,\,|{\bf N}|\leq d\}.$ (7.3)
By (6.3), for $s>0$ and $p\neq 0$, we have
$\max_{\begin{array}[]{c}1\leq k\leq 4\\ {\bf S}_{3}\in[s,\beta_{r}]\end{array}}{\rm Re{\hskip 1.42262pt}}\lambda_{k}({\bf S}_{3})<-\nu(s)<0.$ (7.4)
The tubular domains ${\cal U}^{r}_{d}(a,b)$ with $0<a\leq b<1$ are attracted to $Z^{r}_{+}$ in the following sense.
Lemma 7.1.
Let (7.1) hold, $s\in(0,\beta_{r}/4)$ and $({\bf M}(t),{\bf S}(t))$ be the solution to the averaged equations (4.1) with the initial state $({\bf M}(0),{\bf S}(0))\in D^{r}_{d}:={\cal U}^{r}_{d}(2s,\beta_{r}-2s)$. Then for sufficiently small $d,\mu>0$,
$\max_{t\in[0,p^{-1}]}{\rm dist{\hskip 1.42262pt}}(({\bf M}(t),{\bf S}(t)),Z^{r}_{+})={\cal O}(d).$ (7.5)
Proof.
The Jacobian (6.1) admits a basis of eigenvectors and generalised eigenvectors $v_{k}({\bf S}_{3})$ with ${\bf S}_{3}\in(0,1]$ and $k=1,\dots,5$ corresponding to eigenvalues (6.2).
[…]
$(x_{1},x_{2},x_{3},x_{4},x_{5})\mapsto({\bf M},{\bf S})={\bf Z}^{r}_{2}(x_{5})+\sum_{1}^{4}x_{k}v_{k}(x_{5}),\qquad x_{5}\in(0,\beta_{r}].$ (7.6)
Let us check that the map is nondegenerate at every point $(0,0,0,0,x_{5})$ with $x_{5}\in(0,\beta_{r}]$ which corresponds to $({\bf M},{\bf S})={\bf Z}^{r}_{2}(x_{5})\in Z^{r}_{+}$. Indeed, by (7.2), the last column of the Jacobian $\partial({\bf M},{\bf S})/\partial x$ is the vector ${\partial({\bf M},{\bf S})}/{\partial x_{5}}=(0,0,0,0,1)=v_{5}(x_{5})$. On the other hand for $k\leq 4$, we have ${\partial({\bf M},{\bf S})}/{\partial x_{k}}=v_{k}(x_{5})$. Hence, the columns of the Jacobian are linearly independent. Hence, the coordinates $x({\bf M},{\bf S})$ can be choosen piece-wise continuous in a neighborhood of $Z^{r}_{2}(s,\beta_{r}-s)$. By (7.6) and (5.6), we have $x_{5}={\bf S}_{3}$ for $x_{1}=x_{2}=x_{3}=x_{4}=0$.
Let us fix an $s\in(0,\beta_{r}/4)$ and sufficiently small $d>0$ such that the coordinates $x$ are well defined in ${\cal U}^{r}_{d}(s,\beta_{r}-s)$. Then the distance $d=d(({\bf M},{\bf S}),Z^{r}_{2})$ for $({\bf M},{\bf S})\in{\cal U}^{r}_{d}(s,\beta_{r}-s)$ is equivalent to ${\bf d}={\bf d}(x)=\sqrt{\sum_{1}^{4}x_{k}^{2}}$. For $\delta>0$ let us denote
[…]
Then the coordinates $x$ are well defined in ${\cal T}^{r}_{\delta}(s)$, and the equations (4.1) become
$\left\{\begin{array}[]{rclrcl}\dot{x}_{1}&=&p[\lambda_{1}(x)x_{1}+q_{1}(x)],&\dot{x}_{2}&=&p[\lambda_{2}(x)x_{2}+\varepsilon_{1}(x)x_{1}+q_{2}(x)]\\ \\ \dot{x}_{3}&=&p[\lambda_{3}(x)x_{3}+q_{3}(x)],&\dot{x}_{4}&=&p[\lambda_{4}(x)x_{4}+\varepsilon_{2}(x)x_{3}+q_{4}(x)],\qquad\dot{x}_{5}=pq_{5}(x),\end{array}\right|,$ (7.7)
where $q_{k}(x)={\cal O}({\bf d}^{2})$ and $|\varepsilon_{k}(x)|$ with $k=1,2$ are sufficiently small. By (7.4), the equations (7.7) imply that for solutions $x(t)\in{\cal T}^{r}_{\delta}(s)$ and sufficiently small $\delta>0$,
$\partial_{t}{\bf d}^{2}(x(t))\leq 0.$ (7.8)
Therefore, for any solution $x(t)$ with initial state $x(0)\in{\cal T}^{r}_{\delta}(s)$,
${\bf d}^{2}(x(t))\leq\delta^{2}\quad{\rm until}\quad x_{5}(t)\in[s,\beta_{r}-s].$ (7.9)
This inequality allows us to estimate the exit time $t_{*}$ from ${\cal T}^{r}_{\delta}(s)$ of solutions with initial states $x(0)\in{\cal T}^{r}_{\delta}(2s)$. Namely, in the case $t_{*}<\infty$ and sufficiently small $\delta>0$, the coordinate $x_{5}(t)$ must pass either the segment $[s,2s]$ or $\beta_{r}-2s,\beta_{r}-s]$. In both cases the last equation of (7.7) together with (7.9) imply that
$s/2\leq\int_{0}^{t_{*}}|\dot{x}_{5}(t)|dt\leq pC\int_{0}^{t_{*}}{\bf d}^{2}(x(t))dt\leq pC\delta^{2}\int_{0}^{t_{*}}dt=pC\delta^{2}t_{*}.$ (7.10)
Thus,
$t_{*}>p^{-1}\frac{s}{2C\delta^{2}}.$
Finally, for sufficiently small $\delta>0$ we can choose $s=2C\delta^{2}$. Then (7.9) implies (7.5) for sufficiently small $d>0$. ∎
Remark 7.2.
By (7.4), equations (7.7) with $k\leq 4$ imply an exponential approach to the manifold $Z^{r}_{+}$. Hence, the last equation with $k=5$, gives the exponential decay of $|\dot{x}_{5}|$. This is why the trajectory remains for a long time in a small neighborhood of $Z^{r}_{+}$.
8 Single-frequency asymptotics
In this section, we prove the asymptotics (1.11), (1.12) in the resonance case $\Omega=\omega$. By the relations (3.3), the asymptotics are equivalent to the corresponding asymptotics of solutions $({\mathbb M}(t),{\mathbb S}(t))$ to the interaction dynamics (3.9) which we prove applying the results of the averaging theory [30]. To justify the application, we are going to check suitable properties of the system (3.9).
8.1 KBM vector field
Let us denote $v_{r}=(f_{r},g_{r})$ the vector field (3.10) of the system (3.9). It is easy to check that in the case of almost periodic pumping $A^{e}(t)$ and each $r\!\neq\!0$, for any bounded region ${\cal D}\subset{\mathbb X}$, we have
$\sup_{({\bf M},{\bf S})\in{\cal D}}\frac{1}{T}\Big|\int_{0}^{T}[v_{r}({\bf M},{\bf S},t)-\overline{v}_{r}({\bf M},{\bf S})]dt\Big|\to 0,\qquad T\to\infty.$ (8.1)
Moreover, $v_{r}$ is the Lipschitz continuous vector field in any bounded region ${\cal D}\subset{\mathbb X}$. Hence, $v_{r}$ is a KBM (Krylov–Bogolyubov–Mitropolsky) vector field in any bounded region $\cal D$} according to [30, Definition 4.2.4].
Furthermore, for the quasiperiodic pumping (1.4), the formulas (3.10) implies that
$\delta_{\cal D}(p):=p\sup_{({\bf M},{\bf S})\in{\cal D}}\,\,\sup_{T\in[0,p^{-1}]}\Big|\int_{0}^{T}[v_{r}({\bf M},{\bf S},t)\!-\!\overline{v}_{r}({\bf M},{\bf S})]dt\Big|={\cal O}(p),\,\,\,p\to 0,$ (8.2)
where $\delta_{\cal D}(p)$ is the corresponding order function defined in [30, Lemma 4.6.4].
Remark 8.1.
For more general almost periodic pumping $A^{e}(t)$, the order function can be different [30, Section 4.6].
8.2 The asymptotics in the interaction picture
Here we prove Theorem $1.$
i) The asymptotics (1.11) is equivalent to
$\max_{t\in[0,p^{-1}]}\Big[|{\mathbb M}(t)-{\bf M}|+|{\mathbb S}(t)-{\bf S}|\Big]={\cal O}({p^{1/2}}),\qquad{p\to 0},\quad{p}/\gamma=r,\quad\qquad\qquad\qquad\qquad$ (8.3)
The initial state $({\mathbb M}(0),{\mathbb S}(0))=(M(0),S(0))$ by (3.3). Hence, $({\mathbb M}(0),{\mathbb S}(0))=({\bf M},{\bf S})$. However, $({\bf M},{\bf S})$ is a stationary state for the averaged system (1.9), or equivalently, to (4.7). Therefore, Theorem 4.3.6 of [30] implies (8.3). Indeed, both conditions 1 and 2 of the theorem hold in our case since
a) $v_{r}$ is a KBM-vector field in any bounded region ${\cal D}\subset{\mathbb X}$ with the order function (8.2);
b) solutions $({\mathbb M}(t),{\mathbb S}(t))$ of the system (3.5) with ${p}/\gamma=r$ and small $p>0$ are uniformly bounded by the a priori estimate (A.4).
Hence, (8.3) is proved.
ii) Similarly, the asymptotics (1.12) is equivalent to
$\max_{t\in[0,p^{-1}]}\Big[|{\mathbb M}(t)-{\bf M}_{*}|+|{\mathbb S}(t)-{\bf S}_{*}(t)|\Big]={\cal O}[{p^{1/2}}+d],\qquad p\to 0,\quad{p}/\gamma=r.\qquad$ (8.4)
Define the domain $D^{r}_{d}$ as in Lemma 7.1 and denote by $({\bf M}(t),{\bf S}(t))$ the solution to the averaged system (4.7) with the initial value $({\bf M}(0),{\bf S}(0))=X(0)\in D^{r}_{d}$. Then the same Theorem 4.3.6 of [30] implies that
$\max_{t\in[0,p^{-1}]}|({\mathbb M}(t),{\mathbb S}(t))-({\bf M}(t),{\bf S}(t))|={\cal O}(p^{1/2}),\quad p\to 0,\quad{p}/\gamma=r.$ (8.5)
On the other hand, Lemma 7.1 implies that there exist $({\bf M}_{*}(t),{\bf S}_{*}(t))\in Z^{r}_{+}$ such that
$\max_{t\in[0,p^{-1}]}\Big[|{\bf M}(t)-{\bf M}_{*}(t)|+|{\bf S}(t)-{\bf S}_{*}(t)|\Big]={\cal O}(d).$ (8.6)
Now (8.4) follows since $M_{*}(t)=-{\bf A}^{e}$ by (5.6).
Appendix A The a priori bounds and well-posedness
[…]
Lemma A.1.
There exists a Lyapunov function $V(A,B)$ such that
$a_{1}[A^{2}+B^{2}]\leq V(A,B)\leq a_{1}[A^{2}+B^{2}]\quad{\rm where}\quad a_{1},a_{2}>0,$ (A.1)
and for solutions to (1.1), the function $V(t)=V(A(t),B(t))$ satisfies the inequality
$\dot{V}(t)\leq-\gamma bV(t)+d\frac{p^{2}}{\gamma},\qquad t>0;\quad b,d>0.$ (A.2)
Corollary A.2.
Solving the inequality (A.2), we obtain:
$V(t)\leq V(0)+\frac{d}{b}r^{2},\qquad t\geq 0.$ (A.3)
Hence, for solutions to (1.1) with $p/\gamma=r$, the following bounds hold:
$A^{2}(t)+B^{2}(t)\leq D_{r}(A(0),B(0)),\qquad t\geq 0.$ (A.4)
Now (1.5) follows.
Appendix B The Maxwell–Bloch equations for pure and mixed states
There are various versions of the Maxwell–Bloch equations, see for instance [25] and [2, 3]. In this section we recall introduction of the MBE for pure states and also for mixed states described by density matrix.
2.1 The Maxwell–Bloch equations for pure states
[…]
${\bf A}(x,t)=A(t){\bf X}(x),\quad\psi(x,t)=C_{1}(t)\psi_{1}(x)+C_{2}(t)\psi_{2}(x),\qquad x\in V.$ (B.1)
Here ${\bf A}(x,t)$ denotes the vector potential of the Maxwell field, and ${\bf X}(x)$ is a normalised eigenfunction of the Laplace operator in $V$ under suitable boundary value conditions with an eigenvalue $\Omega^{2}/c^{2}$. By $\psi_{l}$ we denote some normalised eigenfunctions of the Schrödinger operator ${\bf H}:=-\frac{\hbar^{2}}{2{\rm m}}\Delta+e\Phi(x)$ with the corresponding eigenvalues $\hbar\omega_{1}<\hbar\omega_{2}$, where $\Phi(x)$ is the molecular (ion’s) potential. The MBE read as the Hamiltonian system with a dissipation and an external source:
$\frac{1}{c^{2}}\dot{A}(t)=\partial_{B}H,\quad\frac{1}{c^{2}}\dot{B}(t)=-\partial_{A}H-\frac{\gamma}{c^{2}}B;\qquad i\hbar\dot{C}_{l}(t)=\partial_{\overline{C}_{l}}H,\quad l=1,2.$ (B.2)
Here the Hamiltonian is defined as
$H(A,B,C_{1},C_{2},t)={\cal H}(A{\bf X},B{\bf X},C_{1}\varphi_{1}+C_{2}\varphi_{2},t),$
[…]
$H(A,B,C,t)=\frac{1}{2c^{2}}[B^{2}+\Omega^{2}A^{2}]+\hbar\omega_{1}|C_{1}|^{2}+\hbar\omega_{2}|C_{2}|^{2}-\frac{2\varkappa}{c}[A+A^{e}(t)]\,{\rm Im{\hskip 1.42262pt}}[\overline{C}_{1}C_{2}],\quad C=(C_{1},C_{2}).$ (B.3)
Now the Hamilton equations (B.2) become
$\left\{\!\!\!\!\begin{array}[]{rcl}\dot{A}(t)\!\!\!&\!\!\!=\!\!\!&\!\!\!B(t),\quad\dot{B}(t)=-\Omega^{2}A(t)\!-\!\gamma B(t)+cj(t)\\ \\ i\hbar\dot{C}_{1}(t)\!\!\!&\!\!\!=\!\!\!&\!\!\!\hbar\omega_{1}C_{1}(t)+ia(t)\,C_{2}(t),\,\,\,i\hbar\dot{C}_{2}(t)=\hbar\omega_{2}C_{2}(t)-ia(t)\,C_{1}(t)\end{array}\right|,\,\,\,j(t)={2}\varkappa\,{\rm Im{\hskip 1.42262pt}}[\overline{C}_{1}(t)C_{2}(t)].$ (B.4)
The charge conservation $|C_{1}(t)|^{2}+|C_{2}(t)|^{2}=\mathop{\rm const}\nolimits$ follows by differentiation from the last two equations of (B.4) since the function $a(t)$ is real-valued. We consider solutions with $\mathop{\rm const}\nolimits=1$, cf. (1.2):
$|C_{1}(t)|^{2}+|C_{2}(t)|^{2}=1,\qquad t>0.$ (B.5)
2.2 The von Neumann equation for mixed states
The Schrödinger amplitudes $C_{1}(t)$ and $C_{2}(t)$ in (1.1) can be replaced by density matrix $\rho$ which is a nonnegative Hermitian $2\times 2$-matrix (1.2). In particular, for the wave function (B.1), the corresponding density matrix reads as
$\rho(t)=|C(t)\rangle\langle\overline{C}(t)|=\left(\!\!\!\begin{array}[]{cc}|C_{1}(t)|^{2}&C_{1}(t)\overline{C}_{2}(t)\\ \\ C_{2}(t)\overline{C}_{1}(t)&|C_{2}(t)|^{2}\end{array}\!\!\!\right),\qquad C(t)=\left(\!\!\!\begin{array}[]{c}C_{1}(t)\\ C_{2}(t)\end{array}\!\!\!\right).$ (B.6)
The matrix satisfies the trace condition in (1.2) by the charge conservation (B.5). The last line of (B.4) in the vector form reads
$i\hbar\dot{C}(t)=H(t)C(t),\qquad H(t)=\left(\!\!\!\begin{array}[]{cc}\hbar\omega_{1}&ia(t)\\ \\ -ia(t)&\hbar\omega_{2}\end{array}\!\!\!\right)\!\!.$ (B.7)
Accordingly, the density matrix (B.6) satisfies von Neumann equation [32]:
$i\hbar\dot{\rho}(t)=i\hbar[\dot{C}(t)\otimes\overline{C}(t)+C(t)\otimes\dot{\overline{C}}(t)]=[H(t)C(t)]\otimes\overline{C}(t)-C(t)\otimes[H(t)\overline{C}(t)]=[H(t),\rho(t)].$ (B.8)
Now the Maxwell–Bloch system (B.4) is replaced by (1.1).
General density matrix describes an ensemble of molecules with the pure states $C(k,t)=(C_{1}(k,t),C_{2}(k,t))$ satisfying (B.5):
$\rho(t)=\sum_{k}p_{k}C(k,t)\otimes\overline{C}(k,t),\qquad p_{k}\geq 0,\quad\sum_{k}p_{k}=1,$ (B.9)
where $p_{k}$ are the probabilities of the pure states.
Appendix C On possible treatment of the laser action
Here we discuss possible treatment of the laser action relying on the obtained results.
On the smallness of the parameters. Note that the dipole moment $p$ and dissipation coefficient $\gamma$ are very small for many types of lasers. In particular, the dissipation coefficient for the Ruby laser is the electrical conduction of corundum which is $\gamma\sim 10^{-14}$ in the Heaviside–Lorentz units [13, 33, 34]. For the dipole moment typically $|p|\sim 10^{-18}$ according to [26] that agrees with the classical dipole moment $ed/2$, where $d\sim 10^{-8}$cm is the molecular diameter, and $e\sim 10^{-10}$ is the elementary charge in the same units.
On the laser threshold. The asymptotics (1.11) hold for solutions to (1.1) with harmonic initial states $X(0)\in Z^{r}$, and the Lebesgue measure $|Z^{r}|=0$. On the other hand, (1.12) hold for solutions with initial states from an open domain of attraction. Hence, by (1.13), the asymptotics (1.12) appear with a “nonzero probability” in contrast to (1.11). This fact, provisionally, clarifies the existence of a laser threshold to ignite the laser action: the intensity of random pumping must be sufficiently large to bring the solution to the domain of attraction, and then the solution is captured in the domain with the single-frequency asymptotics.
On the laser amplification. The equations (1.1) describe one molecule coupled to the Maxwell field. The limiting amplitudes of the Maxwell field in the asymptotics (1.11), (1.12) do not depend on non-resonance harmonics in the pumping (1.4) with the frequencies $\Omega_{k}\neq\omega$. This means that the dynamics (1.1) acts as a filter, selecting only the resonant harmonics, that itself cannot explain the amplification of the Maxwell field in laser devices. The amplification could be explained by a large number of active molecules, typically $N\sim 10^{20}$, under the traditional assumption that the molecules interact with the Maxwell field but do not interact with each other [28]. In this case, the amplitude of the total Maxwell field is multiplied by $\sqrt{N}\sim 10^{10}$ by the Law of Large Numbers.
This conclusion essentially depends on the fact that the phases of the amplitudes ${\bf M}^{r}$ and $\bf M$$_{*}$ in asymptotics (1.11)–(1.12) for all molecules are approximately uniformly distributed. The uniformity takes place if it holds for phases of ${\bf A}^{e}$. This follows from the asymptotics and Remark 5.2 since $e^{i\phi}e^{-i\Omega t}{\bf M}$ differs from $e^{-i\Omega t}{\bf M}$ by a shift of time.
Self-induced transparency. The harmonic states (5.6) means that the incident wave ${\bf A}^{e}e^{-i\Omega t}$ results in the outgoing wave $M(t)\approx-{\bf A}^{e}e^{-i\Omega t}$ with the same amplitude and frequency, but with the phase jump $\pi$. This phenomenon resembles the self-induced transparency [33].